# Invisible Ears at Your Fingertips: Acoustic Eavesdropping via Mouse Sensors

### v1

@@ I Introduction
The proliferation of low-cost, high-fidelity sensors in consumer devices has greatly improved user experience in common computing tasks. From lower response times to more adaptive workflows, these devices have increased productivity while remaining affordable to the average consumer. The lion’s share of these improvements is found in the category of user input devices, including styli [1, 2], mice [3, 4], and monitors [5, 6]. More specifically, improvements in mice sensor technologies have allowed commercial offerings to operate with a sample rate of 4KHz [7], with a growing selection of products that also support 8KHz [8].
@@
Consumer-grade mice with high-fidelity sensors are already available for under 50 U.S. Dollars [7]. As improvements in process technology and sensor development continue, it is reasonable to expect further price declines, similar to the trend shown in Figure 1. Furthermore, mouse sensors’ resolution and tracking accuracy also follow the same pattern, with steady improvements each year. Ultimately, these developments entail an increased usage of vulnerable mice by consumers, companies, and government entities, expanding the attack surface of potential vulnerabilities in these advanced sensor technologies.
The rise in work-from-home policies has led to the widespread adoption of new technologies and practices, making it more difficult for employers and government institutions to control the physical operating environments of their workforce. While these arrangements often boost employee sentiment and productivity [9], the security implications of work-from-home policies are still being understood [10]. Specifically, attacks exploiting personal peripherals on work computers, such as keyboards [11, 12], microphones [13], styli [14, 15], earphones [16], mechanical hard drives [17], and even USB devices [18], have become increasingly common. Even in relatively secure office environments, the threat posed by these exploits is still significant, especially for unknown or poorly understood attack vectors.
@@ II-A Optical Computer Mice
Modern optical mice employ various methods to provide precise movement tracking under different sensitivity settings. Over the last two decades, optical mice leveraging a high-performance CMOS camera with an onboard Digital Signal Processor (DSP) have become the preferred design choice [19]. Generally, optical sensors enhance reliability and fidelity through the use of self-illumination, typically from an independent diode or an integrated laser. By taking thousands of snapshots of the illuminated surface under the mouse, the DSP can then compare each successive image in order to determine the direction of movement. The rate at which this process happens is determined by the sensor’s frame rate, measured in Frames Per Second (FPS). Each frame is processed via an on-chip correlation algorithm to provide a 2-dimensional displacement to the host computer [20]. The described process can be broken down into two key elements: (i) the imaging sensors and (ii) the image processing and movement detection algorithm.
@@
Imaging Sensors. Rather than relying on expensive Charge-Coupled Device (CCD) sensors, the sensor in an optical mouse is typically a Complementary Metal-Oxide-Semiconductor (CMOS) image sensor collecting up to 30x30 pixels’ worth of data per frame, where each pixel represents the intensity of the reflected light at that point [21]. This basic mini-camera is a critical component for implementing speckle-pattern detection. Some sensor models, such as the PixArt PMW3552, capture data using an 18×18 pixel grid, while others can record up to 30×30 pixels, depending on the manufacturer’s specifications. For visualization purposes, we destructively studied a PixArt PMW3552 sensor in our institutional lab, and the resulting scan is shown in Figure 3. This sensor features an 18×18 CMOS pixel grid and is designed to interface directly via USB. Speckle patterns are random, granular intensity patterns produced when coherent light, such as laser light, is scattered by a rough surface [22]. When an optical mouse is moved across a surface, the speckle pattern on the surface changes smoothly and reliably. The CMOS sensor captures these changes in the speckle pattern frame by frame and processes them to detect movement[23]. These movement detection algorithms allow for the translation of data into corresponding coordinate deltas.
@@
The use of the Fast Fourier Transform (FFT) significantly enhances the efficiency of calculating the correlation function [24]. In the case of image correlation, the FFT is applied to both images ($\vec{I}_{t}$ and $\vec{I}_{t-1}$) to transform them into the frequency domain. This transformation simplifies the calculation of the cross-correlation, which is computationally less intensive in the frequency domain [25]. The equation for cross-correlation in the frequency domain can be expressed as:
@@
The displacement vector $\vec{\Delta p}$, once obtained, is scaled by the DPI setting of the input device to ensure accurate correspondence between physical movement and on-screen cursor movement [26, 27]. This vector is further refined through Kalman filtering, a recursive algorithm enhancing motion tracking by reducing noise in the trajectory of the displacement vector [28, 29]. Finally, this scaled and smoothed vector undergoes quantization in the mouse’s firmware, converting it from a continuous to a discrete form for compatibility with digital systems, thereby maintaining cursor movement accuracy and fluidity, with high-performance mice featuring custom enhancements for input quality.
@@ II-B1 Kernel Space System Calls
Windows operating system also provides a method to retrieve low-level mouse events. By using the LowLevelMouseProc function described in the Windows API, a program can obtain displacement information directly from the operating system [30]. The function yields a data struct of the MSLLHOOKSTRUCT type, containing both the displacement and button information.
@@ II-B2 User Space Applications
Modern operating systems provide a user-space access to input devices. An adversary readily get the data logged by any peripheral (e.g., computer mouse) without necessitating elevated permissions. Many software libraries already provide such access points. Graphical frameworks like Qt[31], GTK[32], and SDL[33] enable real-time mouse data access from unprivileged user-space applications. Similarly, game engines like Unity[34], PyGame [35], and Unreal [36] offer extensive mouse tracking support and built-in networking capabilities, further facilitating an attacker’s efforts.
@@ III Threat Model
Attack Assumptions: We assume that the victim is in a controlled environment, such as a private office or home workspace, where they use a desktop computer equipped with a commercially available optical mouse. The mouse is assumed to have high sensitivity (e.g., a sensor of at least 20,000 DPI) and a polling rate of at least 8 KHz, specifications commonly found in mid-range to high-performance input devices (See Appendix -B). The work surface is assumed to allow vibrations to propagate to the mouse sensor (e.g., a thickness of no more than 3cm), which is typical for wooden or laminate desks with standard thickness. The victim’s computer system is assumed to run an application or process capable of logging pointer movements, such as background services, custom software, or system utilities, which may or may not require elevated privileges. This logging capability is presumed to exist either as a legitimate feature or via unauthorized exploitation discussed in Section III-A. During the attack scenario, the victim is assumed to engage in activities involving sensitive spoken information. While the victim may intermittently use the mouse, the attack leverages instances where the pointer is stationary or experiences minimal motion. The conversation volume is assumed to fall within a range of 60–80dB, which covers typical office or home discussions [37, 38, 39]. The attacker is assumed to have access to mouse movement data via software-level logging, remote exploits, or physical access to the machine. To strengthen the practicality of the threat model, we discuss in Section III-A multiple exploits, including but not limited to compromised applications, unauthorized monitoring software, or vulnerabilities in the input data handling pipeline.
@@ III-A1 Vulnerability Exploit in Graphical Applications
: Software that naturally gathers high-frequency mouse movement data as part of its normal operation can be exploited for the Mic-e-Mouse attack. Images/video editing applications are potential candidates due to their reliance on precise input handling. Examples of such applications include Blender [40], Kdenlive [41], and Krita [42], which utilize mouse data for features like fine-grained control and graphical interface interactions. Many applications with graphical interfaces collect telemetry data, which may include mouse movement logs for diagnostic purposes. While privacy policies typically regulate such data collection, telemetry practices vary widely. In some cases, raw input data could be transmitted to external servers for processing or analysis. A practical attack could involve compromising a telemetry server to gain access to raw mouse movement data. This data, if sufficiently high in resolution, could be analyzed to infer sensitive information, such as speech patterns. A detailed example of this attack and exploit are depicted in Figure 5.
@@ III-A2 Vulnerability Exploit in Open-source Games
: Anti-cheat systems in gaming, such as Vanguard [43], often analyze user input, including mouse movements, to detect tampering or unfair gameplay. In some cases, this analysis involves transmitting input data to remote servers managed by the game developer. While not all anti-cheat systems store raw mouse logs, those that do may create a potential attack vector. However, testing such scenarios is challenging due to the proprietary nature of most closed-source games. For our proof-of-concept attack, we utilize OpenBlok, an open-source version of the classic Tetris video game [44]. Open-source software provides a transparent environment for controlled research and experimentation. Our exploit involves applying a single patch to the game’s source code before compilation. This patch introduces a background thread that captures mouse input data during gameplay and transmits it to a remote server using integrated telemetry. The exploit does not require physical access to the victim’s machine, showcasing its potential applicability in real-world scenarios.
@@ IV-A Microphone Mechanics
Microphones convert sound waves into electrical signals through diaphragm vibrations caused by air pressure changes. The diaphragm’s movement is then transduced into electrical currents via different mechanisms depending on the microphone type [45]. The left side of Figure 6 depicts the internals of a basic microphone. Typically, dynamic microphones generate signals by moving a coil within a magnetic field, whereas condenser microphones alter capacitance by using a vibrating diaphragm as one plate of a capacitor.
@@ IV-A1 How Mice Mimic Microphones
Computer mice are not originally designed to capture sound, but their sensors can detect surface vibrations similar to how a microphone’s diaphragm captures sound waves [20]. When a mouse is stationary, these vibrations can resemble ambient noise picked up by a microphone. An optical mouse’s laser or LED illuminates the surface while the sensor, functioning like a camera, captures thousands of images per second. The mouse’s built-in processor interprets changes between these images to detect movement, which may inadvertently contain sound-induced signals. The unintended microphone-like behavior occurs when the surface vibrations are within the detectable range of the mouse’s sensor. These vibrations cause minute movements of the mouse relative to the surface, which the sensor interprets as movement. With the right signal processing, these signals can be converted back into sound waves, effectively turning the mouse into a rudimentary microphone. Although mice can track these minute vibrations under controlled conditions, real-world usage and surface irregularities can dramatically reduce viability. We thus emphasize that these results represent an upper bound under idealized setups. We show the key similarities between a microphone and the proposed Mic-E-Mouse makeshift microphone in Figure 6. Although conceptually similar to a microphone, the mouse sensor is not designed to respond to air pressure changes. Instead, it passively measures surface displacement, requiring far stronger coupling between acoustic vibrations and the work surface for effective signal capture.
@@ IV-A2 Decoding Vibrations
For a mouse with a given DPI and polling rate, the smallest detectable movement is determined by its resolution. A mouse with a DPI of $D$ can detect movements as small as $1/D$ inches. The polling rate $S$, measured in Hz, is the rate at which the mouse sensor updates the host PC. For transverse waves on a surface with frequency $f$, the mouse sensor will detect these vibrations if their frequency is below the Nyquist frequency, which is half the polling rate. According to the Nyquist-Shannon sampling theorem [46], this is the maximum frequency that can be accurately sampled without aliasing. For a wave with frequency $f$, wavelength $\lambda$ on the surface can be calculated if we know the speed of the wave on the surface $\nu$:
@@ IV-B Preliminary Study
We perform a preliminary study to confirm that the mouse can detect frequencies within the range of human speech—typically between 200 and 2 KHz [47, 48]. We generate a step-wise sweep from 0 to 16 kHz to observe any aliased components at frequencies above the Nyquist limit (4 kHz), thus testing how the mouse sensor responds to out-of-band signals. The sweep is generated using a synthetic tone and sweep audio sample which consists of four consecutive segments: (i) an initial 200Hz tone lasting 5 seconds, (ii) a linear sweep from 0 to 1 KHz over 4 seconds, (iii) a linear sweep from 0 to 16 KHz over 4 seconds, and (iv) a 400Hz tone lasting 5 seconds. The start and end tones serve as a synchronization token since we observed a strong response during our initial experiments. Additionally, we cross-reference the spectral distribution of the logged mouse signals, the true data signals, and a collection of background noise in order to establish a statistical model of the noise profile for further processing during the signal processing stage of our pipeline. Figure 8 presents: (a) spectral densities of ground truth speech, ambient sensor noise, and speech collected from the mouse; (b) the ground truth synthetic sweep; (c) the sweep collected from the mouse; and (d) a speech signal collected from the mouse. In Figure 8(a), the mouse-captured signal exhibits two prominent peaks that align in frequency with the ground truth but differ in amplitude, indicating partial resonance unique to the sensor. Meanwhile, panel (c) highlights aliased harmonics that provide further evidence of how effectively (or poorly) the mouse replicates subtle audio content. The mouse sensor shows a non-Gaussian noise profile with a power peak around 1 kHz. Figures 8b and 8c demonstrate a ’tone and sweep’ test, with the mouse-collected data mirroring the ground truth pattern, confirming the presence of a side-channel. In Figure 8d, speech signals appear faint but distinct below 500 Hz, indicating the need for classical and neural filtering techniques.

### v2

@@ I Introduction
The proliferation of low-cost, high-fidelity sensors in consumer devices has greatly improved user experience in common computing tasks. From lower response times to more adaptive workflows, these devices have increased productivity while remaining affordable to the average consumer. The lion’s share of these improvements is found in the category of user input devices, including styli [ngstyli, kodakstyli], mice [faresmouse, xiangmouse], and monitors [amirmonitor, marcumonitor]. More specifically, improvements in mice sensor technologies have allowed commercial offerings to operate with a sample rate of 4KHz [darmosharkm3], with a growing selection of products that also support 8KHz [razerviper].
@@
Consumer-grade mice with high-fidelity sensors are already available for under 50 U.S. Dollars [darmosharkm3]. As improvements in process technology and sensor development continue, it is reasonable to expect further price declines, similar to the trend shown in Figure 1. Furthermore, mouse sensors’ resolution and tracking accuracy also follow the same pattern, with steady improvements each year. Ultimately, these developments entail an increased usage of vulnerable mice by consumers, companies, and government entities, expanding the attack surface of potential vulnerabilities in these advanced sensor technologies.
The rise in work-from-home policies has led to the widespread adoption of new technologies and practices, making it more difficult for employers and government institutions to control the physical operating environments of their workforce. While these arrangements often boost employee sentiment and productivity [klinewfh], the security implications of work-from-home policies are still being understood [borkewfh]. Specifically, attacks exploiting personal peripherals on work computers, such as keyboards [zhukeyboard, asonovkeyboard], microphones [genkinlend], styli [farrukhpencil, liumaghacker], earphones [liecho], mechanical hard drives [kwonghdd], and even USB devices [dumitruusb], have become increasingly common. Even in relatively secure office environments, the threat posed by these exploits is still significant, especially for unknown or poorly understood attack vectors.
@@ II-A Optical Computer Mice
Modern optical mice employ various methods to provide precise movement tracking under different sensitivity settings. Over the last two decades, optical mice leveraging a high-performance CMOS camera with an onboard Digital Signal Processor (DSP) have become the preferred design choice [palacinoptical]. Generally, optical sensors enhance reliability and fidelity through the use of self-illumination, typically from an independent diode or an integrated laser. By taking thousands of snapshots of the illuminated surface under the mouse, the DSP can then compare each successive image in order to determine the direction of movement. The rate at which this process happens is determined by the sensor’s frame rate, measured in Frames Per Second (FPS). Each frame is processed via an on-chip correlation algorithm to provide a 2-dimensional displacement to the host computer [gordondsp]. The described process can be broken down into two key elements: (i) the imaging sensors and (ii) the image processing and movement detection algorithm.
@@
Imaging Sensors. Rather than relying on expensive Charge-Coupled Device (CCD) sensors, the sensor in an optical mouse is typically a Complementary Metal-Oxide-Semiconductor (CMOS) image sensor collecting up to 30x30 pixels’ worth of data per frame, where each pixel represents the intensity of the reflected light at that point [cmossensor]. This basic mini-camera is a critical component for implementing speckle-pattern detection. Some sensor models, such as the PixArt PMW3552, capture data using an 18×18 pixel grid, while others can record up to 30×30 pixels, depending on the manufacturer’s specifications. For visualization purposes, we destructively studied a PixArt PMW3552 sensor in our institutional lab, and the resulting scan is shown in Figure 3. This sensor features an 18×18 CMOS pixel grid and is designed to interface directly via USB. Speckle patterns are random, granular intensity patterns produced when coherent light, such as laser light, is scattered by a rough surface [stachowiakspeckle]. When an optical mouse is moved across a surface, the speckle pattern on the surface changes smoothly and reliably. The CMOS sensor captures these changes in the speckle pattern frame by frame and processes them to detect movement[cmosspeckle]. These movement detection algorithms allow for the translation of data into corresponding coordinate deltas.
@@
The use of the Fast Fourier Transform (FFT) significantly enhances the efficiency of calculating the correlation function [fft]. In the case of image correlation, the FFT is applied to both images ($\vec{I}_{t}$ and $\vec{I}_{t-1}$) to transform them into the frequency domain. This transformation simplifies the calculation of the cross-correlation, which is computationally less intensive in the frequency domain [ft]. The equation for cross-correlation in the frequency domain can be expressed as:
@@
The displacement vector $\vec{\Delta p}$, once obtained, is scaled by the DPI setting of the input device to ensure accurate correspondence between physical movement and on-screen cursor movement [boudmouse, pangmouse]. This vector is further refined through Kalman filtering, a recursive algorithm enhancing motion tracking by reducing noise in the trajectory of the displacement vector [kalmanfilter, likalman]. Finally, this scaled and smoothed vector undergoes quantization in the mouse’s firmware, converting it from a continuous to a discrete form for compatibility with digital systems, thereby maintaining cursor movement accuracy and fluidity, with high-performance mice featuring custom enhancements for input quality.
@@ II-B1 Kernel Space System Calls
Windows operating system also provides a method to retrieve low-level mouse events. By using the LowLevelMouseProc function described in the Windows API, a program can obtain displacement information directly from the operating system [winapimouse]. The function yields a data struct of the MSLLHOOKSTRUCT type, containing both the displacement and button information.
@@ II-B2 User Space Applications
Modern operating systems provide a user-space access to input devices. An adversary readily get the data logged by any peripheral (e.g., computer mouse) without necessitating elevated permissions. Many software libraries already provide such access points. Graphical frameworks like Qt[Qt], GTK[GTK], and SDL[SDL] enable real-time mouse data access from unprivileged user-space applications. Similarly, game engines like Unity[unity], PyGame [pygame], and Unreal [unreal] offer extensive mouse tracking support and built-in networking capabilities, further facilitating an attacker’s efforts.
@@ III Threat Model
Attack Assumptions: We assume that the victim is in a controlled environment, such as a private office or home workspace, where they use a desktop computer equipped with a commercially available optical mouse. The mouse is assumed to have high sensitivity (e.g., a sensor of at least 20,000 DPI) and a polling rate of at least 8 KHz, specifications commonly found in mid-range to high-performance input devices (See Appendix -B). The work surface is assumed to allow vibrations to propagate to the mouse sensor (e.g., a thickness of no more than 3cm), which is typical for wooden or laminate desks with standard thickness. The victim’s computer system is assumed to run an application or process capable of logging pointer movements, such as background services, custom software, or system utilities, which may or may not require elevated privileges. This logging capability is presumed to exist either as a legitimate feature or via unauthorized exploitation discussed in Section III-A. During the attack scenario, the victim is assumed to engage in activities involving sensitive spoken information. While the victim may intermittently use the mouse, the attack leverages instances where the pointer is stationary or experiences minimal motion. The conversation volume is assumed to fall within a range of 60–80dB, which covers typical office or home discussions [nassilamp, LidarPhone, gyrophone]. The attacker is assumed to have access to mouse movement data via software-level logging, remote exploits, or physical access to the machine. To strengthen the practicality of the threat model, we discuss in Section III-A multiple exploits, including but not limited to compromised applications, unauthorized monitoring software, or vulnerabilities in the input data handling pipeline.
@@ III-A1 Vulnerability Exploit in Graphical Applications
: Software that naturally gathers high-frequency mouse movement data as part of its normal operation can be exploited for the Mic-e-Mouse attack. Images/video editing applications are potential candidates due to their reliance on precise input handling. Examples of such applications include Blender [swblender], Kdenlive [swkdenlive], and Krita [swkrita], which utilize mouse data for features like fine-grained control and graphical interface interactions. Many applications with graphical interfaces collect telemetry data, which may include mouse movement logs for diagnostic purposes. While privacy policies typically regulate such data collection, telemetry practices vary widely. In some cases, raw input data could be transmitted to external servers for processing or analysis. A practical attack could involve compromising a telemetry server to gain access to raw mouse movement data. This data, if sufficiently high in resolution, could be analyzed to infer sensitive information, such as speech patterns. A detailed example of this attack and exploit are depicted in Figure 5.
@@ III-A2 Vulnerability Exploit in Open-source Games
: Anti-cheat systems in gaming, such as Vanguard [vanguard], often analyze user input, including mouse movements, to detect tampering or unfair gameplay. In some cases, this analysis involves transmitting input data to remote servers managed by the game developer. While not all anti-cheat systems store raw mouse logs, those that do may create a potential attack vector. However, testing such scenarios is challenging due to the proprietary nature of most closed-source games. For our proof-of-concept attack, we utilize OpenBlok, an open-source version of the classic Tetris video game [openblok]. Open-source software provides a transparent environment for controlled research and experimentation. Our exploit involves applying a single patch to the game’s source code before compilation. This patch introduces a background thread that captures mouse input data during gameplay and transmits it to a remote server using integrated telemetry. The exploit does not require physical access to the victim’s machine, showcasing its potential applicability in real-world scenarios.
@@ IV-A Microphone Mechanics
Microphones convert sound waves into electrical signals through diaphragm vibrations caused by air pressure changes. The diaphragm’s movement is then transduced into electrical currents via different mechanisms depending on the microphone type [hoppermic]. The left side of Figure 6 depicts the internals of a basic microphone. Typically, dynamic microphones generate signals by moving a coil within a magnetic field, whereas condenser microphones alter capacitance by using a vibrating diaphragm as one plate of a capacitor.
@@ IV-A1 How Mice Mimic Microphones
Computer mice are not originally designed to capture sound, but their sensors can detect surface vibrations similar to how a microphone’s diaphragm captures sound waves [gordondsp]. When a mouse is stationary, these vibrations can resemble ambient noise picked up by a microphone. An optical mouse’s laser or LED illuminates the surface while the sensor, functioning like a camera, captures thousands of images per second. The mouse’s built-in processor interprets changes between these images to detect movement, which may inadvertently contain sound-induced signals. The unintended microphone-like behavior occurs when the surface vibrations are within the detectable range of the mouse’s sensor. These vibrations cause minute movements of the mouse relative to the surface, which the sensor interprets as movement. With the right signal processing, these signals can be converted back into sound waves, effectively turning the mouse into a rudimentary microphone. Although mice can track these minute vibrations under controlled conditions, real-world usage and surface irregularities can dramatically reduce viability. We thus emphasize that these results represent an upper bound under idealized setups. We show the key similarities between a microphone and the proposed Mic-E-Mouse makeshift microphone in Figure 6. Although conceptually similar to a microphone, the mouse sensor is not designed to respond to air pressure changes. Instead, it passively measures surface displacement, requiring far stronger coupling between acoustic vibrations and the work surface for effective signal capture.
@@ IV-A2 Decoding Vibrations
For a mouse with a given DPI and polling rate, the smallest detectable movement is determined by its resolution. A mouse with a DPI of $D$ can detect movements as small as $1/D$ inches. The polling rate $S$, measured in Hz, is the rate at which the mouse sensor updates the host PC. For transverse waves on a surface with frequency $f$, the mouse sensor will detect these vibrations if their frequency is below the Nyquist frequency, which is half the polling rate. According to the Nyquist-Shannon sampling theorem [shannonfs], this is the maximum frequency that can be accurately sampled without aliasing. For a wave with frequency $f$, wavelength $\lambda$ on the surface can be calculated if we know the speed of the wave on the surface $\nu$:
@@ IV-B Preliminary Study
We perform a preliminary study to confirm that the mouse can detect frequencies within the range of human speech—typically between 200 and 2 KHz [bishopfreq, speechbanana]. We generate a step-wise sweep from 0 to 16 kHz to observe any aliased components at frequencies above the Nyquist limit (4 kHz), thus testing how the mouse sensor responds to out-of-band signals. The sweep is generated using a synthetic tone and sweep audio sample which consists of four consecutive segments: (i) an initial 200Hz tone lasting 5 seconds, (ii) a linear sweep from 0 to 1 KHz over 4 seconds, (iii) a linear sweep from 0 to 16 KHz over 4 seconds, and (iv) a 400Hz tone lasting 5 seconds. The start and end tones serve as a synchronization token since we observed a strong response during our initial experiments. Additionally, we cross-reference the spectral distribution of the logged mouse signals, the true data signals, and a collection of background noise in order to establish a statistical model of the noise profile for further processing during the signal processing stage of our pipeline. Figure 8 presents: (a) spectral densities of ground truth speech, ambient sensor noise, and speech collected from the mouse; (b) the ground truth synthetic sweep; (c) the sweep collected from the mouse; and (d) a speech signal collected from the mouse. In Figure 8(a), the mouse-captured signal exhibits two prominent peaks that align in frequency with the ground truth but differ in amplitude, indicating partial resonance unique to the sensor. Meanwhile, panel (c) highlights aliased harmonics that provide further evidence of how effectively (or poorly) the mouse replicates subtle audio content. The mouse sensor shows a non-Gaussian noise profile with a power peak around 1 kHz. Figures 8b and 8c demonstrate a ’tone and sweep’ test, with the mouse-collected data mirroring the ground truth pattern, confirming the presence of a side-channel. In Figure 8d, speech signals appear faint but distinct below 500 Hz, indicating the need for classical and neural filtering techniques.